OpenStudy (anonymous):

The length of a rectangle is 5 centimeters less than twice its width.The perimeter of the rectangle is 26 cm. What are the dimensions of the rectangle?

OpenStudy (anonymous):

what is the formula for perimeter?

OpenStudy (anonymous):

p = l+w

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

2l+2w

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

26 = 2(width) + 2(width - 5) 26= 2(width+width-5) 13 = width + width - 5 18 = width + width 9= width 9-5 = 4 = length 4+4+9+9 = 26

OpenStudy (anonymous):

difference of 5!

OpenStudy (anonymous):



OpenStudy (anonymous):

x=7...

OpenStudy (anonymous):

raphael, stop harrassing people on other questions, its not very nice, and then deleting your hurtful words!

OpenStudy (emilylauren):

do you still need help?

OpenStudy (emilylauren):

x=6

OpenStudy (emilylauren):

to find perimeter we need length and width. Width=x Length=2x-5

OpenStudy (emilylauren):

2L+2W=26

OpenStudy (emilylauren):

so 2(x)+2(2x-5)=26 2x+4x-10=26

OpenStudy (emilylauren):

6x=36 x=6 Hope I helped. Please make sure you understand the problem as well as know the answer :)

OpenStudy (emilylauren):

width is 6 length is 7